The marks of 500 candidates in an examination are normally distributed 

with mean of 45 marks and a standard deviation of 20 marks. 

a) Given that the pass mark is 41, estimate the number of candidates 

who passed the examination.

b) If 5% of the candidates obtained a distinction by scoring x marks or 

more, estimate the value of x.


1
Expert's answer
2021-05-20T14:58:12-0400

Let "Y=" the mark in an examination: "Y\sim N(\mu, \sigma^2)"

Given "\mu=45, \sigma=20."


a)


"P(Y\geq41)=1-P(Y<41)"

"=1-P(Z<\dfrac{41-45}{20})=1-P(Z<-0.2)"

"\approx0.579260"

"500(0.579260)=289.63"

289 candidates passed the examination.


b)


"P(Y\geq x)=1-P(Y<x)"

"=1-P(Z<\dfrac{x-45}{20})=0.05"

"P(Z<\dfrac{x-45}{20})=0.95"

"\dfrac{x-45}{20}=1.6449"


"x=78"